Part A What is the pH of a buffer prepared by adding 0.506 mol of the weak acid HA to 0.608 mol of NaA in 2.00 L of so...

Part A

What is the pH of a buffer prepared by adding 0.506 mol of the weak acid HA to 0.608 mol of NaA in 2.00 L of solution? The dissociation constant Ka of HA is 5.66×10−7.

Express the pH numerically to three decimal places.

pH =

SubmitHintsMy AnswersGive UpReview Part

Part B

What is the pH after 0.150 mol of HCl is added to the buffer from Part A? Assume no volume change on the addition of the acid.

Express the pH numerically to three decimal places.

pH =

SubmitHintsMy AnswersGive UpReview Part

Part C

What is the pH after 0.195 mol of NaOH is added to the buffer from Part A? Assume no volume change on the addition of the base.

A - C

Express the pH numerically to three decimal places.

pH =

SubmitHintsMy AnswersGive UpReview Part
Answer #1

pH = bka + log [Naat [HA] pH = -log Ka + dog 0.608 0.506 bH= 6:326 HA. 0:506 A + H+ = initial: 0.608 - adding at: - 0.150 fin+ OK A 0.608 + 4,0 HA initial: 0.506 adding on: - finally: 0.311 0.195 0 0.803 - bu = b ka + dog [A] [CHA] pH = 7 - 2095.66